# Integrating Random Regret Minimization-Based Discrete Choice Models with Mixed Integer Linear Programming for Revenue Optimization

### v2

@@ 3.2.5 Capacitated supply MILP model
A capacitated version of the previously presented model has been depicted in this part. The term ”capacitated” is used to emphasize that both resources and the available quantity of an alternative are limited.
@@
Constraints 19, 20, 30 ensure that once an alternative is not in the choice set of a customer, it will never be chosen by the customer.
Constraints 33, 34, 35 ensure that it is possible to have $y_{inr}=1$ when all the customers before the $n_{i}$ have been served, and there exist enough quantity of alternative $i$ for the $n_{i}$ customer.
The above constraints do not offer any other available alternative to the customer. Constraints associated with discounted regret, 25,26, 27, 28, and 29 help find the available alternatives with the least anticipated regret in the choice set of a customer. Constraints 30 , 31, 32 address the issue of picking up the alternative with the least regret while it has enough quantity to be allocated. The rest of the constraints are the same as in the previous model. Finally, constraints 19, 20, 33, 34, 35, 36, 37, 39 were taken from pacheco2016new .
@@ 4 Computational experiments
In this section, we employ IBM ILOG CPLEX Optimization to execute the models and assess their performance in solving the numerical experiments.
In design of the experiments, we included three alternatives in the choice set for each customer. One of them represents the no-buy option, while the other two alternatives are left unlabeled. Regret arises due to the price attribute and unobserved factors. We did not consider the socio-economic characteristics of the customers in this experiment, i.e., $ER_{ijnr}=0,\ \forall i,j,n,r$. Specifically, price is the only attribute for both alternatives. Customers are homogeneous in their preferences, so $\beta_{np}=-1,\ \forall n.$ Also, $v_{onr}$ and $v_{nr}$ are positive random draws from $Gumbel(0,1)$ distribution. Price has six levels ranging from $1$ to $4.5$, with $pm=0.5$. The bounds are set to be arbitrarily large enough numbers. As for the RUM-based MILP modeled by pacheco2017integrating , the overall utility is solely associated with the price attribute and the unobserved utility, which is a random draw from $Gumbel(0,1)$.
Firstly, we solved both capacitated and uncapacitated RRM-based MILP models to optimality. We considered $r=4$ draws for each customer in both models. Table 1 summarizes the performance of the models. To check if capacity constraints are active, examine the instance with $11$ customers. There are five units of each of alternatives. Thus, there is only the opt-out choice for one of the customers, which shows itself in the gap between the uncapacitated and capacitated solution.
@@
Also, it is important to note that despite having the same number of customers and equal capacity for the alternatives (except in the case of 11 customers), the uncapacitated model was solved more quickly. This could be attributed to the fact that the capacitated version involves additional constraints to adhere to capacity restrictions, leading to delays in solving the model. Additionally, we conclude that the opt-out option was never chosen given the increasing revenue trend while the capacity of each of alternatives has increased by one unit.
Given that the customers and alternatives are identical, the number of feasible solutions and iterations tends to be high due to possible symmetries. The models involve a substantial number of variables and constants, even for a relatively small number of customers and draws, as depicted in Table 2. The combinatorial nature of these models necessitates the application of heuristic methods to address computational challenges.
@@
Table 3 holds particular significance. In the RRM-based model, higher price levels have been selected, in contrast to the RUM-based model where lower price levels are favored. This discrepancy highlights the impact of different choice behavior theories on results, emphasizing the importance of utilizing these theories appropriately in demand and supply studies. Applying RUM-based models where customers are regret minimizers can lead to huge profit losses.
@@
The structure of the RRM-based model, which compares alternatives pairwise, introduces many symmetries in this particular experiment, thereby increasing the number of iterations, as shown in Table 4. Additionally, there are more variables and constraints in the RRM-based MILP compared to its RUM-based counterpart signaling potential application of heauristics and more sophisticated models for RRM-based MILPs to tackle large size problems.
@@
In conclusion, this experiment demonstrated the feasibility of our models and underscored the significance of studying different consumer choice behaviors, as they can impact revenue, pricing, and capacity allocation policies. Our results demonstrate the superiority of our RRM-based MILP in revenue maximization when customers are regret minimizers.
@@ 5 Conclusion
In conclusion, cutting-edge Revenue Management (RM) techniques, adept at handling the intricate dynamics of supply and demand, leverage Discrete Choice Models (DCMs) as superior demand estimators rooted in consumer purchasing behavior. These models are seamlessly integrated into Mixed-Integer Linear Programming (MILP) frameworks, drawing inspiration from Random Utility Maximization (RUM) theory and the more recent Random Regret Minimization (RRM) theory. Notably, RUM-based DCMs have been the go-to choice in the RM landscape.
This study contributes by formulating MILP models tailored for both capacitated and uncapacitated supply scenarios, with a specific focus on customers as regret minimizers based on RRM theory. The results not only validate the viability of these models but also underscore the substantial impact of varied customer behavior interpretations on revenue, pricing, and customer choices.
Future research endeavors may entail tackling large-scale datasets to solve RUM-based and RRM-based MILP models, shedding light on their real-world applicability. In the face of potential computational challenges, especially with the heightened complexity of RRM-based MILP models, adopting Benders-type heuristics, column generation, or meta heuristics emerge as a pragmatic solution.
Furthermore, recognizing that both RUM-based and RRM-based MILPs yield extreme values for prices, utility, and regret, the development of bi-objective MILPs is proposed. Introducing a second objective, aimed at maximizing utility or minimizing regret, could offer deeper insights. Exploring Pareto optimal frontiers in such models promises nuanced information about customer satisfaction, empowering suppliers with a more informed approach to decision-making in Revenue Management.

### v3

@@ 3.2.5 Capacitated supply MILP model
A capacitated version of the previously presented model has been depicted in this part. The term ”capacitated” is used to emphasize that the available quantity of an alternative is limited.
@@
Constraints 22, 23, 33 are incorporated into the model to ensure that once an alternative is not in the choice set of a customer, it will never be allocated to that customer. When $y_{in}=0$, variables $y_{inr}$ are imposed to be zero (23) and consequently, $w_{inr}=0$ due to constraints (33).
Constraints 36, 37, and 38 ensure that the capacity of the alternatives is respected. Here is how it works: Since in this version of the model, we assume that all the alternatives are available to all customers unless capacity is limited if $y_{inr}=0$, it means that there is not enough quantity of alternative $i$ for the customer $n$ at scenario $r$. Thus, it will not be available for the customer $n+1$, and consequently, $y_{i(n+1)r}=0$ due to constraints 36. Then, constraints 37 and 38 will imply that there is not enough quantity for the alternative $i$, since constraint 37 will be $\sum_{m=1}^{n-1}w_{imr}\leq n-1$, meaning that at most $n-1$ unit of alternative $i$ has been allocated to earlier customers (customers 1 to $n-1$), and constraint 38 will be $c_{i}\leq\sum_{m=1}^{n-1}w_{imr}$, meaning that all $c_{i}$ quantity has been allocated to customers before customer $n$. If $y_{inr}=1$, constraints 36 imply that the customer $n+1$ may still purchase alternative $i$. Also, constraints 37 will imply that $\sum_{m=1}^{n-1}w_{imr}\leq c_{i}-1$, i.e., at most $c_{i}-1$ unit of alternative $i$ has been allocated to customers earlier than customer $n$, and still, there is some capacity of alternative $i$ to be allocated to customer $n$.
Constraints 24, 25, 26 and 27 have been explained previously. However, the key difference between the capacitated model and the uncapacitated model is that there might not be an available quantity of an alternative in the capacitated model. Hence, the anticipated regret should be computed based on the available alternatives for the customer. Constraints associated with discounted regret, 28,29, 30, 31, and 32 help find the available alternatives with the least anticipated regret in the choice set of a customer. Therefore, we propose $z_{ijnr}$, discounted regret variables, $z_{ijnr}=R_{ijnr}$ while $y_{jnr}=1$ and $z_{ijnr}=0$ when $y_{jnr}=0$. In other words, when the competing alternative is not available due to capacity limitations, there is no regret, and the anticipated regret should be computed among available alternatives. The idea is similar to that of RUM-based MILP developed by [2]. However, in our model, since the regret is a result of the pairwise comparison of alternatives, it is important to check if both alternatives are available. Constraints 28,29, and 30 work as follows:
If $y_{jnr}=1$, then
$z_{ijnr}\leq RR_{ijnr}+ER_{ijnr}$
$RR_{ijnr}+ER_{ijnr}\leq z_{ijnr}$
implying that
$z_{ijnr}=RR_{ijnr}+ER_{ijnr}.$
Otherwise,
$z_{ijnr}\leq RR_{ijnr}+ER_{ijnr}$
$RR_{ijnr}+ER_{ijnr}-M\leq z_{ijnr}$
We know that $M$ is big enough, so
$RR_{ijnr}+ER_{ijnr}-M\leq 0.$
Thus, $z_{ijnr}$ can take any value in $[RR_{ijnr}+ER_{ijnr}-M,RR_{ijnr}+ER_{ijnr}]$. But, we want this variable to be zero. Constraints 30 ensure that while $y_{jnr}=0$, then $z_{ijnr}=0$. They will be inactive once $y_{jnr}=1$ since $0\leq z_{ijnr}\leq M$ always holds. Consequently, to avoid the effect of the not-available alternative $j$ on the overall anticipated regret of the alternative $i$, we replace the constraint set
$RR_{ijnr}=\sum_{j\neq i,j\in C_{n}}RR_{ijnr}+ER_{ijnr},\quad\forall i\in C_{n},n,r,$
with the constraint set
$RR_{ijnr}=\sum_{j\neq i,j\in C_{n}}z_{ijnr},\quad\forall i\in C_{n},n,r.$
Constraints 33, 34, 35 address the issue of picking up the alternative with the least regret while it has enough quantity to be allocated.
The rest of the constraints are the same as in the previous model. Finally, constraints 22, 23, 36, 37, 38, 39, 40, 42 were taken from [2].
@@ 4 Computational experiments
In our experiments, we included three alternatives in the choice set for each customer. One of them represents the no-buy option, while the other two alternatives are left unlabeled. Regret arises due to the price attribute and unobserved factors. We did not consider the socio-economic characteristics of the customers in this experiment, i.e., $ER_{ijnr}=0,\ \forall i,j,n,r$. Specifically, price is the only attribute for both alternatives. Customers are homogeneous in their preferences, so $\beta_{np}=-1,\ \forall n.$ Also, $v_{onr}$ and $v_{nr}$ are positive random draws from $Gumbel(0,1)$ distribution. Price has six levels ranging from $1$ to $4.5$, with $pm=0.5$. The bounds are set to be arbitrarily large enough numbers. As for the RUM-based MILP modeled by [15], the overall utility is solely associated with the price attribute and the unobserved utility, which is a random draw from $Gumbel(0,1)$.
Firstly, we solved both capacitated and uncapacitated RRM-based MILP models to optimality. We considered $r=4$ draws for each customer in both models, i.e., we considered 4 scenarios for each customer.
4.1 Numerical results
We employ IBM ILOG CPLEX Optimization solver to execute the models and assess their performance in solving the numerical experiments.
The Table 1 summarizes the performance of both capacitated and uncapacitated models. The maximum revenue achievable in our experiments could be obtained by multiplying the number of customers by the highest price level, i.e., \$ $\$\ 4.5$. Thus, all the results in the table imply that once there is a sufficient quantity of alternatives for each customer, maximum revenue is attained. In other words, regret stemming from the price of the alternatives is negligible compared to regret when choosing the opt-out option. This could be partly because we have only included the regret arising from the price of the alternatives. To verify if capacity constraints are active, consider the instance with 11 customers. In this case, there are five units of each alternative available. Consequently, only the opt-out choice is available for one of the customers, which is reflected in the gap between the uncapacitated and capacitated solutions.
It is also important to note that despite having the same number of customers and equal capacity for the alternatives (except in the case of 11 customers), the uncapacitated model was solved more quickly. This could be attributed to the fact that the capacitated version involves additional constraints to adhere to capacity restrictions, leading to delays in solving the model.
@@ 4.1 Numerical results
In Table 2, the number of constraints and variables for both models shows that the capacitated model contains more variables and constraints. This is due to the introduction of capacity-related constraints with additional variables $z_{ijnr}$, $y_{in}$, and $y_{inr}$. The additional variables and constraints, stemming from the combinatorial nature of these models, generally require more time for the model to be solved, as described in Table 2. Also, since we did not label the alternatives, and customers are identical, there could be a lot of symmetries in the model. This means that the solution method of the model distinguishes between the same solutions, thereby increasing solving time. The branch and cut method was used by the CPLEX solver to find the solutions. Indeed, the ’nodes’ column in Table 2 demonstrates the number of nodes on the search tree of the branch and cut algorithm. A large number of nodes for the capacitated version shows that the algorithm assessed the same solutions many times due to the symmetry of the model, most possibly because of the capacity-related constraints, as these numbers are fewer in the uncapacitated model. Additionally, ’iterations’ show how many times the branch and cut algorithm was used to solve the model. Certainly, if the number of nodes is high, the number of iterations will be high as well, since each node inspected by the algorithm implies additional runs of the algorithm.
All in all, the addition of capacity-related constraints has increased the solution time, which was expected due to the combinatorial nature of the models.
@@
In Table 3, we have shown the results of the experiment for the uncapacitated version of the RRM-based and RUM-based MILP models. These models represent different theories of customer decision-making: the RRM-based model and the RUM-based model taken from [2].
In the RRM-based model, which is rooted in behavioral economics, customers are assumed to make decisions based on minimizing regret. This means they choose options that minimize the disappointment or dissatisfaction they might feel after making a decision. On the other hand, the RUM-based model is based on traditional economic theory and assumes that customers make decisions to maximize their utility or satisfaction.
The results presented in Table 3 indicate that in the RRM-based model, higher price levels tend to be selected by customers. This suggests that customers in this model are willing to pay more to avoid regret. Conversely, in the RUM-based model, customers favor lower price levels. This indicates that customers in this model are more focused on maximizing their utility by seeking lower prices.
This discrepancy highlights the impact of different choice behavior theories on results, emphasizing the importance of utilizing these theories appropriately in demand and supply studies. Applying RUM-based models where customers are regret minimizers can lead to significant revenue losses in the real world, as evidenced by the observed differences in revenue (the ’Gap’ column) between the two models.
Note that in our experiments, we have not assumed the true behavior of the customers, i.e., we have not pre-assumed that customers are regret minimizers or utility maximizers in order to obtain unbiased results. Our goal is to demonstrate the potential implications of different customer decision-making theories, such as RRM and RUM, on the outcomes of our models.
The observed ’Gap’ in the solutions of RRM-based and RUM-based models underscores the importance of understanding and accurately modeling customer behavior. If customers are indeed regret minimizers, as suggested by the results of the RRM-based model, then implementing such models could benefit both suppliers and customers. Suppliers can optimize revenue by offering price levels that minimize customer regret, while customers benefit from choices that align with their decision-making preferences.
Conversely, applying an RRM-based model in scenarios where customers are utility maximizers, as indicated by the results of the RUM-based model, may not be favorable for customers. In such cases, customers may perceive the offerings as less attractive or may experience dissatisfaction due to higher prices, potentially leading to revenue losses for suppliers.
The fact that the number of customers affects the number of variables, thereby increasing the size of the search tree in the Branch and Cut algorithm, results in longer solving time. This effect, however, is more pronounced for the RRM-based MILP as it contains more variables than its RUM-based counterpart. However, neither model provides insight into the regret or utility levels of customers for different price levels. Such information can significantly help suppliers devise more effective marketing and production strategies. For instance, developing bi-objective MILPs enables the generation of Pareto frontiers, which depict the trade-offs between prices and revenue for various levels of customer regret or utility. This allows suppliers to make informed decisions that optimize both revenue and customer satisfaction.
@@
In Table 4, the structural properties of RUM-based and RRM-based MILPS for several instances are presented. As the number of customers increases, both models exhibit an increase in the number of variables and constraints. However, the RRM-based MILP demonstrates a larger number of variables and constraints compared to its RUM-based counterpart. This disparity arises from the additional constraints and variables required by the RRM theory, which need to be linearized and integrated into the MILP model.
Furthermore, Table 4 reveals that the solver required significantly more iterations to solve the RRM-based MILP compared to its RUM-based counterpart. This indicates that for large-sized problems with numerous customers, the RRM-based model will require a substantial amount of time to achieve optimality. Consequently, operations research techniques such as metaheuristic methods and decomposition techniques should be developed so that the RRM-based MILP can obtain approximate or exact solutions in a timely manner.
@@ 5 Conclusion
In conclusion, cutting-edge Revenue Management (RM) techniques, adept at handling the intricate dynamics of supply and demand, leverage Discrete Choice Models (DCMs) as superior demand estimators rooted in consumer purchasing behavior. These models are seamlessly integrated into Mixed-Integer Linear Programming (MILP) frameworks, drawing inspiration from Random Utility Maximization (RUM) theory and the more recent Random Regret Minimization (RRM) theory. Notably, RUM-based DCMs have been the go-to choice in the RM landscape. However, numerous pieces of research show that RRM-based DCMs are superior to RUM-based DCMs in scenarios in which customers tend to minimize regret.
This study contributes by formulating novel MILP models tailored for both capacitated and uncapacitated supply scenarios, with a specific focus on customers as regret minimizers based on RRM theory.
Particularly, the anticipated regret of all alternatives was computed through the pairwise comparison of all alternatives using MILP model constraints. Additionally, price-related regret included price level variables to compute regret for different price levels. Simultaneously, price level variables were integrated into the objective function of the MILP models to maximize revenue. By doing so, demand and supply were linked to both maximize revenue while accounting for customer regret minimization behavior.
The results not only validate the viability of these models but also underscore the substantial impact of varied customer behavior interpretations on revenue, pricing, and customer choices.
uture research endeavors may involve addressing large-scale datasets to solve RUM-based and RRM-based MILP models, shedding light on their real-world applicability. Specifically, the structures of MILP models indicate a significant number of variables and constraints, even for small instances. Given the potential computational challenges, particularly with the heightened complexity of RRM-based MILP models, adopting Benders-type heuristics, column generation, or meta-heuristics emerges as a pragmatic solution.
Furthermore, acknowledging that both RUM-based and RRM-based MILPs often result in extreme values for prices, utility, and regret, the development of bi-objective MILPs is proposed. Introducing a second objective aimed at maximizing utility or minimizing regret could provide deeper insights. Exploring Pareto optimal frontiers, i.e., investigating the trade-offs between revenue or prices and utility or regret, in such models promises nuanced information about customer satisfaction, empowering suppliers with a more informed approach to decision-making in Revenue Management.